what is the perimeter of a rectangle that has an area of 12,150 square feet? the ratio of length to width is 3:2.
Answer by dabanfield(803)

what is the perimeter of a rectangle that has an area of 12,150 square feet? the ratio of length to width is 3:2.
Let the two sides of the rectangle be 3x and 2x respectively. Then we have:
3x*2x = 12150
6x^2 = 12150
x^2 = 2025
x = 45
The two sides then are 2*45 = 90 and 3*45 = 135